Charter Communications, Inc. provides both cable TV (X) and high speed Internet service (Y) through their Spectrum brand. Suppose the cost of providing different levels of the service is given below, where the cost of providing different levels of the services is given by C (X,Y):

C (0, 40) = 50

C (0, 80) = 90

C (100, 0) = 110

C (200, 0) = 250

C (100, 40) = 150

C (200, 80) = 320

Does Charter have economies of scale in the production of either service? Does it have economies of scope in the production of both services together? Explain
Answer #1

Economies of scale

Economies of scale in the production of goods or services is said to e achieved, when they are produced on a large scale with more units of goods or service with low input cost.If economies of scale is achieved ,economic growth is said to be realised.

Economies of scale in the production of high speed Internet service (Y)

Units Total Cost Cost per Unit
40 50 1.25
80 90 1.125

Cost Economies of Scale - 1.25 1.125 - 40 8 Output

Economies of scale in the production of cable TV(X)

Units Total Cost Cost per units
100 110 1.1
200 250 1.25

Avg Cost Diseconomies k of I scale 7 1.25 1. 27 1.151 100 200 Outon

It shows that economies of scale is obtained in the case of high speed internet service and not in the case of cable TV.

Economies of scope

In the case of, C (100, 40) = 150

Total cost of producing service together=150

Cost of producing goods separately; Cost of producing cable TV service =100 x 1.1=110

Cost of producing high speed internet service =40 x 1.25=50

Total cost =100+50=160

Here economies of scope is obtained since total cost of producing services together is less than total Cost of producing goods separately

In the case of, C (200, 80) = 320

Total cost of producing service together=320

Cost of producing goods separately; Cost of producing cable TV service =200x1.25=250

Cost of producing high speed internet service =80 x 1.125=90

Total cost =90+250=340

Here economies of scope is obtained since total cost of producing services together is less than total Cost of producing goods separately